The following data are pulled from a recent Walsh Manufacturing annual report.

Assets

Raw material inventory                   $130,000

Work-in-process inventory               $60,000

Finished goods inventory                $310,000

Property, plant & equipment          $500,000

Other assets                                   $200,000

Total assets                                   $1,200,000

Condensed Income Statement

Revenue                                        $2,000,000

Cost of goods sold                           $600,000

Other expenses                             $1,000,000

Net income                                      $400,000

Calculate:

(a) Percent invested in inventory

(b) Inventory turnover

(c) Weeks of supply.

Please provide detail work on each calculation.
Answer #1

A) Percent Invested in Inventory can be computed by dividing the total inventory by total assets.

Raw Material Inventory = $130,000

WIP Inventory = $60,000

Finished Goods Inventory = $310,000

Total Inventory = 130,000 + 60,000 + 310,000

= $500,000

Total assets = $1,200,000

Percent Invested in inventory = Total Inventory / Total Assets

= 500,000 / 1,200,000

= 0.4167 = 41.67%

B) Inventory Turnover = Cost of Goods Sold / Average Inventory

Also,

Inventory Turnover = Sales / Total Inventory

= 2,000,000 / 5,00,000

= 4 Times

C) Weeks of Supply =Number of working weeks in a year / Inventory Turnover ratio

= 52 / 4

= 13 Weeks